OpenStudy (anonymous):

Can someone please help me this is literally making no sense. I've tried to learn this but I suck at math. Solve the inequality and graph the solution. 8|x+3/4|<2

OpenStudy (jdoe0001):

first off, keep in mind that |something| really means +(something) AND -(something) so you really end up with 2 answers, or 2 values for "x" \({\bf 8\left|x+\cfrac{3}{4}\right|< 2\\ \textit{dividing each side by 8}\\ \left|x+\cfrac{3}{4}\right|<\cfrac{2}{8} \implies \left|x+\cfrac{3}{4}\right|<\cfrac{1}{4}\\ -----------------\\ \left|x+\cfrac{3}{4}\right|<\cfrac{1}{4} \implies } \begin{cases} +\left(x+\cfrac{3}{4}\right)<\cfrac{1}{4} \implies &x+\cfrac{3}{4}<\cfrac{1}{4}\\ -\left(x+\cfrac{3}{4}\right)<\cfrac{1}{4} \implies &-x-\cfrac{3}{4}<\cfrac{1}{4} \end{cases} \)

OpenStudy (anonymous):

omg thank you so much! T_T

OpenStudy (jdoe0001):

yw